# Fedor V. Fomin

### v2

@@ Abstract
The classic theorem of Gallai and Milgram (1960) asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$. The proof of Gallai–Milgram’s theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths.
We prove the following algorithmic extension of Gallai–Milgram’s theorem for undirected graphs: determining whether an undirected graph can be covered by fewer than $\alpha(G)-k$ vertex-disjoint paths is fixed-parameter tractable ($\mathsf{FPT}$) when parameterized by $k$. More precisely, we provide an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time
$2^{k^{O(k^{4})}}\cdot n^{O(1)},$
and outputs a path cover $\mathcal{P}$ of $G$. Furthermore, it:
  • either correctly reports that $\mathcal{P}$ is a minimum-size path cover,
  • or outputs, together with $\mathcal{P}$, an independent set of size $\lvert\mathcal{P}\rvert+k$ certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
A key subroutine in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path. This result is of independent interest—prior to our work, no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most 3. Moreover, the techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment. We show that all these problems are $\mathsf{FPT}$ when parameterized by the independence number of the graph.
Notably, the independence number parameterization, which describes graph’s density, departs from the typical flow of research in parameterized complexity, which focuses on parameters describing graph’s sparsity, like treewidth or vertex cover.
@@ 1 Introduction
The Gallai–Milgram Theorem, published by Tibor Gallai and Arthur Milgram in 1960 [GM60], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s Theorem [Dil50], concerning the covering of graphs by simpler substructures. In its original form, the theorem states that every (directed or undirected) graph $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the size of a maximum independent set in $G$.
Beyond its immediate implications, the Gallai–Milgram Theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence. This theme can be traced back to the seminal work of Rédei [Réd34], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path. Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [NW71] and Chvátal–Erdős [CE72], and have continued in later works (see, e.g., [AFG83, BBVR89, Jac87, Koi94, BAH88, BT07, HJ90, Tho01, FS09]).
In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among the Gallai–Milgram theorem, Hamiltonicity, and the independence number.
1.1 Our contribution.
Known constructive proofs of the Gallai–Milgram theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths. Our following theorem presents a significant generalization of this algorithmic result for undirected graphs: for an integer parameter $k$, deciding whether a graph $G$ contains a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf FPT}}$, parameterized by $k$.
@@
Theorem 1 aligns with a broader trend in parameterized algorithms that seeks algorithmic analogues of classic theorems in extremal graph theory. Examples include algorithmic extensions of Poljak–Turzík’s theorem [MPSS14], Dirac’s theorem [FGSS22a], and Hajnal–Szemerédi’s theorem by Gan, Han, and Hu [GHH23]. As with Theorem 1, each of these results provides an $\operatorname{{\sf FPT}}$ algorithm parameterized “above” or “beyond” a known combinatorial bound. A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [DF99] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22]. Consequently, we cannot simply require $\alpha(G)$ to be part of the input, which makes the existence of an $\operatorname{{\sf FPT}}$ algorithm parameterized below $\alpha(G)$ all the more surprising.
As a special case of Theorem 1, let $k=\alpha(G)-1$. In this scenario, the problem reduces to determining whether all vertices of $G$ can be covered by a single path—i.e., whether $G$ has a Hamiltonian path. Consequently, Theorem 1 implies that deciding whether $G$ contains a Hamiltonian path is $\operatorname{{\sf FPT}}$ when parameterized by $\alpha(G)$. The construction of such an $\operatorname{{\sf FPT}}$ algorithm for Hamiltonicity is a key component of the proof of Theorem 1. Moreover, the techniques developed to achieve this result extend well beyond Hamiltonicity and apply to a broader class of problems.
Similar to Theorem 1, the situation concerning parameterization by $\alpha(G)$ is unusual: Most algorithmic findings in structural parameterization require that computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$. For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ parameterized by the treewidth of a graph due to the following reasoning. First, for an input integer $k$, there are (either exact or approximation) algorithms that either correctly decide that the treewidth of the input graph is more than $k$ or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [Bod96, Kor21, RS86]. Then, the tree decomposition is used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$. In contrast, this time we deal with the parameter that does not admit an $\operatorname{{\sf FPT}}$ approximation [LRSW22, KK22].
Our second theorem provides an $\operatorname{{\sf FPT}}$ algorithm for computing topological minors in a graph, parameterized by its independence number.
@@
Let $H$, $G$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of $M\setminus T$ have degree $2$. We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to a graph obtained from $M$ by dissolving in $M$ all vertices of $M\setminus T$. Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent. We further say that for the injective mapping $f:V(H)\to V(G)$ induced by the isomorphism above, $(M,f)$ is a topological minor embedding (TM-embedding) of $H$ in $G$. The size of a TM-model $(M,T)$ and of a TM-embedding $(M,f)$ is the number of vertices in $M$.
@@
The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$ parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$. The algorithm, for a given parameter $k$, either computes a TM-embedding of maximum size in $\operatorname{{\sf FPT}}$ time with respect to $k$, or correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$. Such algorithms are what Raghavan and Spinrad referred to as robust [RS03]. More precisely,
@@
There is an algorithm that, given graphs $H,G$, a list assignment $L:V(H)\to 2^{V(G)}$ and an integer parameter $k\geq 1$, in time $2^{|H|^{\mathcal{O}(k)}}\cdot 2^{k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$
@@
To our knowledge, neither list TM-embeddings nor the Maximum List TM-Embedding problem have been studied so far. However, this problem is a natural generalization of several vital problems in graph algorithms. In particular, Theorem 2 implies the following corollaries.
@@
  • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$. If this path contains $|V(G)|$ vertices, this is a Hamiltonian path. For $H=C_{3}$, we have the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$. By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$.
@@
  • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths. By Gallai–Milgram’s theorem, if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$. We can solve the problem by taking $H$ to be disjoint unions of at most $p$ paths $P_{1}$ and $P_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$. Thus, Theorem 2 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem. Of course, it also follows from Theorem 1. For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
@@
  • Consider the following generalization of Path Cover. In the $-\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$. The task is to connect these pairs by disjoint paths that altogether traverse all vertices of the graph. For $\ell=1$ this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path. $-\ell$-Linkage could be encoded as List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$. Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$. Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$. By Theorem 2, $-\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$. By playing with different graphs $H$ in the List TM-embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
@@
  • Another interesting scenario is the maximization variant of the $T$-Cycle problem. In the $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals. The task is to decide whether there is a cycle passing through all terminals [BHT12, Kaw08, Wah13] (in [BHT12] authors look for the minimum-length cycle). We can reduce the maximization variant to finding a list TM-embedding of maximum size by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution. For each guessed ordering, we apply Theorem 2 for $H$ being a cycle on $|T|$ vertices such that each vertex of $H$ is assigned to precisely one terminal vertex, and the cyclic ordering of $H$ corresponds to the guess. This brings us to the $f(\alpha(G)+|T|)\cdot|G|^{\mathcal{O}(1)}$ algorithm computing a cycle of maximum length containing all $T$ terminals. Similar arguments hold for a more general colored variant of the problem. In this variant, we are given a graph $G$, and some of the vertices of $G$ are assigned colors from $1$ to $L$. Then, the task is to find a longest cycle containing the maximum number of colors [FGK+23]. Note that an optimal cycle can contain many vertices of the same color. By guessing the colors participating in a solution cycle $C$ and the ordering of their first appearance in some cyclic ordering of $C$, we construct a cycle $H$ on at most $L$ vertices. Every vertex of $H$ is assigned the list of vertices of $G$ corresponding to the guessed color. This brings us to the algorithm of running time $f(\alpha(G)+L)\cdot|G|^{\mathcal{O}(1)}$.

### v3

# Path Cover, Hamiltonicity, and Independence Number: An FPT Perspective

@@ Abstract
The classic theorem of Gallai and Milgram (1960) generalizes several fundamental results in Graph Theory, such as Dilworth’s theorem on posets and Kőnig’s theorem on matchings in bipartite graphs. The theorem asserts that for every graph $G$, the vertex set of $G$ can be partitioned into at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the maximum size of an independent set in $G$. The proof of the Gallai–Milgram theorem is constructive and yields a polynomial-time algorithm that computes a covering of $G$ by at most $\alpha(G)$ vertex-disjoint paths. While the Gallai–Milgram theorem is tight—there are graphs where one really needs $\alpha(G)$ paths, not fewer, to cover the vertex set of GG—it was not known prior to our work whether deciding if a graph $G$ could be covered by fewer than $\alpha(G)$ vertex-disjoint paths can be done in polynomial time.
We resolve this question by proving the following algorithmic extension of the Gallai–Milgram theorem for undirected graphs: There is an algorithm that, for an $n$-vertex graph $G$ and an integer parameter $k\geq 1$, runs in time $2^{2^{\mathcal{O}(k^{4}\log{k})}}\cdot n^{\mathcal{O}(1)}$ and outputs a path cover $\mathcal{P}$ of $G$ together with
  • a correct conclusion that $\mathcal{P}$ is a minimum-size path cover, or
  • an independent set of size $\lvert\mathcal{P}\rvert+k$, certifying that $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
Thus, for $k\in\mathcal{O}((\log\log n)^{\frac{1}{4}-\varepsilon})$ our algorithm runs in polynomial time and either computes a minimum-size path cover of $G$, or finds a path cover of size at most $\alpha(G)-k$. We find the existence of such an algorithm quite surprising for the following reason. The problems of computing a path cover and a maximum independent set are both notoriously hard, yet our algorithm either solves one of them or provides meaningful information about the other.
The proof of our algorithmic extension of the Gallai–Milgram theorem is non-trivial and builds on several novel algorithmic ideas. One of the key subroutines in our algorithm is an $\mathsf{FPT}$ algorithm, parameterized by $\alpha(G)$, for deciding whether $G$ contains a Hamiltonian path. This result is of independent interest—prior to our work no polynomial-time algorithm for deciding Hamiltonicity was known even for graphs with independence number at most three. Moreover, the algorithmic techniques we develop apply to a wide array of problems in undirected graphs, including Hamiltonian Cycle, Path Cover, Largest Linkage, and Topological Minor Containment. We show that all these problems are $\mathsf{FPT}$ when parameterized by the independence number of the graph.
Notably, the independence-number parameterization departs from the typical direction of research in parameterized complexity. First, $\alpha(G)$ measures a graph’s density, whereas most prior work in the area focuses on parameters describing sparsity, such as treewidth or vertex cover. Second, most structural parameters studied in parameterized complexity can be computed exactly or well-approximated in polynomial or even $\mathsf{FPT}$ time, whereas computing $\alpha(G)$ is notoriously difficult from almost any computational perspective. The fact that it can nevertheless serve as the basis for efficient parameterization is particularly striking.
@@ 1 Introduction
Gallai–Milgram’s theorem, published by Tibor Gallai and Arthur Milgram in 1960 [45], is a cornerstone of structural graph theory and can be viewed as a far-reaching generalization of several classical results, such as Dilworth’s and Kőnig’s theorems [29, 68], concerning the covering of graphs. In its original form, the theorem states that every (directed or undirected) graph $G$ can be covered by at most $\alpha(G)$ vertex-disjoint paths, where $\alpha(G)$ is the size of a maximum independent set in $G$.
Beyond its immediate implications, Gallai–Milgram’s theorem belongs to a broader line of research in graph theory focused on connections between Hamiltonicity and independence. This theme can be traced back to the seminal work of Rédei [85], who showed that every tournament (a directed graph with $\alpha(G)=1$) contains a directed Hamiltonian path. Further investigations into the relationships among stability number, connectivity, and Hamiltonicity began with Nash-Williams [82] and Chvátal–Erdős [19], and have continued in later works (see, e.g., [3, 5, 58, 67, 6, 8, 57, 88, 43]).
In this paper, we present a novel algorithmic perspective on this classic theme in graph theory: the interplay among Gallai–Milgram’s theorem, Hamiltonicity, and the independence number.
1.1 Our contribution
Known constructive proofs of Gallai–Milgram’s theorem can be used to construct in polynomial time a covering of the vertex set of a graph $G$ by at most $\alpha(G)$ vertex-disjoint paths. Our following theorem presents a significant generalization of this algorithmic result for undirected graphs.
@@ 1.1 Our contribution
Loosely speaking, the theorem means that computing a path cover below Gallai–Milgram’s bound, that is, a path cover of size at most $\alpha(G)-k$, is $\operatorname{{\sf FPT}}$ when parameterized by $k$. However, formally, such a statement is not correct: A trivial reduction shows that deciding whether $G$ has a path cover of size at most $\alpha(G)-k$ is $\operatorname{{\sf W}}[1]$-hard. Indeed, for an $n$-vertex input graph $G$, construct a graph $G^{\prime}$ by adding a clique of size $n$ whose vertices are made universal to $G$. Then $\alpha(G^{\prime})=\alpha(G)$, and $G^{\prime}$ is trivially Hamiltonian. Thus, for $k^{\prime}=k-1$, the graph $G^{\prime}$ can be covered by at most $\alpha(G^{\prime})-k^{\prime}$ paths if and only if $\alpha(G)\geq k$. Since the Independent Set problem is $\operatorname{{\sf W}}[1]$-complete [30] when parameterized by the solution size, the existence of an $\operatorname{{\sf FPT}}$ algorithm in $k$ for deciding whether $G$ has a path cover of size at most $\alpha(G)-k$ would contradict the basic complexity assumptions of parameterized complexity theory.
Of course, in situations where the independence number of a graph can be efficiently computed, for example when the input graph is perfect, Theorem 1 implies that the path cover below Gallai–Milgram is $\operatorname{{\sf FPT}}$. The interesting feature of Theorem 1 is that it circumvents the intractability of computing $\alpha(G)$—the algorithm does not need to know $\alpha(G)$ at all. This is achieved by allowing the computation of an optimal path cover rather than concluding that $G$ has no path cover of size at most $\alpha(G)-k$. Another notable feature of the algorithm is that it always outputs a path cover which is either optimal or of size at most $\alpha(G)-k$. It is somewhat ironic that the “formal” intractability of the path cover below Gallai–Milgram arises only from the requirement of identifying no-instances, that is, situations where $\alpha(G)-k$ is smaller than the size of a minimum path cover.
@@
Theorem 1 aligns with a broader trend in parameterized algorithms that seeks algorithmic analogues of classic theorems in extremal graph theory. Examples include algorithmic extensions of Poljak–Turzík’s theorem [80], Dirac’s theorem [37, 60], and Hajnal–Szemerédi’s theorem [46]. As with Theorem 1, each of these results provides an $\operatorname{{\sf FPT}}$ algorithm parameterized “above” or “beyond” a known combinatorial bound. A distinguishing feature of Theorem 1, however, is that in previous work the relevant combinatorial bounds (e.g., the maximum vertex degree) were efficiently computable, whereas determining whether $\alpha(G)\geq k$ is $\operatorname{{\sf W}}[1]$-hard with respect to $k$ [30] and likely does not admit an $\operatorname{{\sf FPT}}$ approximation [71, 63].
One of the key components in the proof of Theorem 1 is an $\operatorname{{\sf FPT}}$ algorithm, parameterized by $\alpha(G)$, for deciding the Hamiltonicity of $G$. Moreover, the techniques developed to obtain this result extend well beyond Hamiltonicity and apply to a broader class of problems. Our second main result, Theorem 2, provides an $\operatorname{{\sf FPT}}$ algorithm for computing various topological embeddings in a graph, parameterized by its independence number. While only a restricted case of this result is needed in the proof of Theorem 1, we believe that the general approach is of independent interest, and therefore present it in full generality.
@@
Let $G$, $H$ be graphs, let $M$ be a subgraph of $G$ and $T\subseteq V(M)$ be a vertex set such that all vertices of MTM-T have degree two in $M$. We say that $(M,T)$ is a topological minor model (TM-model) of $H$ in $G$ if $H$ is isomorphic to the graph obtained from $M$ by dissolving in $M$ all vertices of MTM-T. Here, dissolving a vertex of degree two means deleting it while making its neighbors adjacent. We further say that for the injective mapping $f:V(H)\to V(G)$ induced by the isomorphism above, $(M,f)$ is a topological minor embedding (TM-embedding) of $H$ in $G$. The size of a TM-model $(M,T)$ and of a TM-embedding $(M,f)$ is the number of vertices in $M$.
@@
The following theorem establishes that computing the list TM-embedding of maximum size is $\operatorname{{\sf FPT}}$, that is, solving the Maximum List TM-Embedding problem, parameterized by the size of the maximum independent set in $G$ and the size of the graph $H$. The algorithm, for a given parameter $k$, either computes a TM-embedding of maximum size in $\operatorname{{\sf FPT}}$ time with respect to $k$, or correctly reports that the input graph does not fall within the appropriate domain, meaning it has an independence number greater than $k$. Such algorithms are what Raghavan and Spinrad refer to as robust [83]. More precisely,
@@
There is an algorithm that, given graphs $H,G$, a list assignment $L:V(H)\to 2^{V(G)}$ and an integer parameter $k\geq 1$, in time $2^{|H|^{\mathcal{O}(k)}\cdot k^{\mathcal{O}(k^{2})}}\cdot|G|^{\mathcal{O}(1)}$
@@
Similar to Theorem 1, the question concerning parameterization by $\alpha(G)$ deviates from the standard setup used in structural parameterized complexity, where most algorithmic results rely on the fact that the problem of computing or approximating the corresponding graph structural parameter is in $\operatorname{{\sf FPT}}$. For example, deciding whether a graph contains a Hamiltonian cycle is $\operatorname{{\sf FPT}}$ when parameterized by the treewidth of the graph, for the following reason. Given an input integer $k$, there exist algorithms (either exact or approximate) that either correctly decide that the treewidth of the input graph is greater than $k$, or construct a tree decomposition of width $\mathcal{O}(k)$ in time $f(k)\cdot n^{\mathcal{O}(1)}$ [12, 69, 86]. The resulting tree decomposition can then be used to decide whether the graph is Hamiltonian in time $g(k)\cdot n^{\mathcal{O}(1)}$. In contrast, in our setting, the parameter under consideration does not admit an $\operatorname{{\sf FPT}}$ approximation [71, 63], which necessitates a different approach.
Theorem 2 implies that Maximum List TM-Embedding is $\operatorname{{\sf FPT}}$ with respect to the combined parameter $|H|+\alpha(G)$. For any integer $k>\alpha(G)$, the algorithm of Theorem 2 either outputs a maximum-size list TM-embedding or correctly determines that no such embedding exists. The fact that $\alpha(G)$ cannot be computed efficiently can be circumvented by the following trick. To find a list TM-embedding of $H$ in $G$, we run the algorithm of Theorem 2 for increasing values of the parameter $k$, starting with $k=1$, until the algorithm either finds an embedding or concludes that no appropriate embedding exists.
@@
To our knowledge, neither list TM-embeddings nor the Maximum List TM-Embedding problem have been studied so far. However, these problems are natural generalizations of several vital problems in graph algorithms. In particular, Theorem 2 implies the following corollaries.
@@
  • When graph $H$ consists of a path $P_{2}$ on two vertices whose lists are all vertices of $V(G)$, then this is the problem of finding a longest path in $G$. If this path contains $|V(G)|$ vertices, this is a Hamiltonian path. For $H=C_{3}$, we have the problem of finding a longest cycle in $G$, the generalization of computing a Hamiltonian cycle in $G$. By Theorem 2, these tasks could be performed in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ which implies that the problem of finding a longest path/cycle is $\operatorname{{\sf FPT}}$ parameterized by $\alpha(k)$.
  • A graph $G$ is Hamilton-connected (also known as Hamiltonian connected) if every two vertices of $G$ are connected by a Hamiltonian path; see, e.g., [16, p. 61]. Checking whether a graph is Hamilton-connected thus reduces to computing, for each pair $u,v\in V(G)$, a maximum-size TM-embedding of $H=P_{2}$ with lists $\{u\}$ and $\{v\}$, which can be done in time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ by Theorem 2.
@@
  • In Path Cover, the task is to cover all vertices of $G$ by at most $p$ vertex-disjoint paths. By Gallai–Milgram’s theorem, if $p\geq\alpha(G)$ then a covering by at most $p$ paths always exists and we can assume that $p<\alpha(G)$. We can solve the problem by taking $H$ to be disjoint unions of at most $p$ disjoint copies of $K_{2}$ assigning to each vertex of $H$ the list containing all vertices of $V(G)$. Thus, Theorem 2 yields an algorithm of running time $f(\alpha(G))\cdot|G|^{\mathcal{O}(1)}$ solving this problem. Of course, it also follows from Theorem 1. For Cycle Cover, that is when the graph $H$ consists of at most $p$ disjoint copies of $C_{3}$, we have the running time $f(\alpha(G)+p)\cdot|G|^{\mathcal{O}(1)}$.
@@
  • Consider the following generalization of Path Cover. In the Hamiltonian-$\ell$-Linkage problem, we are given $\ell$ pairs of vertices $(s_{1},t_{1}),\dots,(s_{\ell},t_{\ell})$. The task is to connect these pairs by disjoint paths that altogether traverse all the vertices of the graph. For $\ell=1$, this is the problem of deciding whether a graph for selected vertices $s$ and $t$ contains a Hamiltonian $s,t$-path. Hamiltonian-$\ell$-Linkage could be encoded as Maximum List TM-embedding with $H$ consisting of $\ell$ disjoint copies of $P_{2}$. Let the vertices of these $P_{2}$’s be $\{x_{i},y_{i}\}$, $i\in\{1,\dots,\ell\}$. Then the list mapping $L$ assigns $\{s_{i}\}$ to $x_{i}$ and $\{t_{i}\}$ to $y_{i}$. By Theorem 2, Hamiltonian-$\ell$-Linkage is solvable in time $f(\alpha(G)+\ell)\cdot|G|^{\mathcal{O}(1)}$. By playing with different graphs $H$ in Maximum List TM-Embedding and Theorem 2, one easily obtains algorithms computing spanning subgraphs with various properties like a spanning tree with at most $\ell$ leaves.
@@
  • Another interesting scenario is the maximization variant of the $T$-Cycle problem. In the $T$-Cycle problem, we are given a graph $G$ and a set $T\subseteq V(G)$ of terminals. The task is to decide whether there is a cycle passing through all terminals [11, 64, 89] (in [11] authors look for the minimum-length cycle). We can reduce the maximization variant to finding a list TM-embedding of maximum size by enumerating all $|T|!$ orderings in which terminal vertices occur in an optimal solution. For each guessed ordering, we apply Theorem 2 for $H$ being a cycle on $|T|$ vertices such that each vertex of $H$ is assigned to precisely one terminal vertex, and the cyclic ordering of $H$ corresponds to the guess. This brings us to the $f(\alpha(G)+|T|)\cdot|G|^{\mathcal{O}(1)}$ algorithm computing a cycle of maximum length containing all terminals in $T$. Similar arguments hold for a more general colored variant of the problem. In this variant, we are given a graph $G$, and some of the vertices of $G$ are assigned colors from $1$ to $L$. Then, the task is to find a longest cycle containing the maximum number of colors [33]. Note that an optimal cycle can contain many vertices of the same color. By guessing the colors participating in a solution cycle $C$ and the ordering of their first appearance in some cyclic ordering of $C$, we construct a cycle $H$ on at most $L$ vertices. Every vertex of $H$ is assigned the list of vertices of $G$ corresponding to the guessed color. This brings us to the algorithm of running time $f(\alpha(G)+L)\cdot|G|^{\mathcal{O}(1)}$.
We remark that the result of Theorem 2 is tight in the sense that Maximum List TM-Embedding is trivially $\operatorname{{\sf Para-NP}}$-hard when parameterized by either $\alpha(G)$ or $|H|$ only. The problem is $\operatorname{{\sf NP}}$-hard for $H=K_{2}$ because of the $\operatorname{{\sf NP}}$-completeness of Hamiltonian Path [47]. Furthermore, observe that finding a path with $n$ vertices in a bipartite $n$-vertex graph $G$ is equivalent to finding the complement of the $n$-vertex path in the complement of $G$ which is cobipartite. Since Hamiltonian Path is $\operatorname{{\sf NP}}$-complete on bipartite graphs [47] and the independence number of a cobipartite graph is at most two, Maximum List TM-Embedding is $\operatorname{{\sf NP}}$-hard for graphs $G$ with $\alpha(G)\leq 2$ when $H$ is a part of the input.